OpenStudy (anonymous):

Simplify each radical expression. -5(square root sign) 112

OpenStudy (anonymous):

112=2*2*2*2*7 so sqrt 112= 4*sqrt7

OpenStudy (anonymous):

\[4\sqrt{7}\]

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

Okay. Thank you hun :)

OpenStudy (anonymous):

np

OpenStudy (anonymous):

if u need to do anything like this just write the number up in multiplication: like this 112=2*2*2*2*7

OpenStudy (anonymous):

Okay :)